and 2) Solve for the general solution of the order system dx 22 (6) Z 12t dzz =(
Answer #1

2-3)+ 1 lety & // x = 20 -202 + 1 x = gdz =x-nä 24,-28+) -24 912 Amilay equation is; 24+24 - 2 + 1 - m u2 + H²=1 mum = 0 mhtu 9() = ezta et -1+ (test) thett luett Say(t) = G+2t+Cs (1+ed) - 2net - 1 - - z (+) = (citet tett luet th 2 9, (2, (3,ly ar